# Error Correction Code Transformer: From Non-Unified to Unified

### v3

@@ 3.1 Shared Unified Attention
The vanilla self-attention mechanism in Transformer architectures establishes dense correlations among all input features within a sequence, resulting in a computational complexity of $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$, where $N$ is the sequence length and $d_{k}$ is the embedding dimension. This approach, while effective for capturing long-range dependencies, is computationally intensive and overlooks the intrinsic relationships across different samples, particularly in channel decoding tasks involving diverse linear block codes. Moreover, prior studies [30] suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution. This sparsity implies that a low-rank approximation could effectively capture the dominant correlations, reducing redundancy without compromising performance.
@@
To investigate this hypothesis, we trained an ECCT model [24] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including Polar(32,16), Polar(64,32), Polar(128,86), LDPC(49,24), LDPC(121,60), LDPC(121,80), BCH(31,16), BCH(63,36), and BCH(127,120). We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the six-layer Transformer encoder. As illustrated in Figure 5, the rank ranges from 12 to 181, with over 90% of values below 69, consistently indicating a low-rank structure across all layers and code types. This observation suggests that the high-dimensional self-attention matrix can be approximated with a lower-rank representation, offering a pathway to reduce computational overhead.
@@
Furthermore, the conventional multi-head self-attention mechanism computes independent attention scores for each head using distinct weight matrices $\textbf{{W}}_{i}^{Q}$, $\textbf{{W}}_{i}^{K}$, and $\textbf{{W}}_{i}^{V}$. However, we hypothesize that the attention distributions across heads may exhibit high similarity, particularly in channel decoding, where the sparse structure of the parity-check matrix $\mathbf{H}$ constrains bit interactions. To quantify this, we employed the Jensen-Shannon Divergence (JSD) [31], a symmetric measure of similarity between probability distributions, defined as:
$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right]$ (9)
@@
$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right)$ (10)
@@
$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right)$ (11)
By eliminating the projections $\textbf{{W}}^{\textit{Q}}$ and $\textbf{{W}}^{\textit{K}}$, the UAM reduces the computational complexity from $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$ to $\mathcal{O}(N\cdot d_{l}\cdot d_{k})$, where $d_{l}\ll N$. The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to foster a cohesive decoding architecture. This design not only reduces the parameter count but also enables the model to learn shared features across diverse code types, facilitating a code-agnostic framework.
@@ 3.2 Sparse Masked Unified Attention
$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right)$ (12)
@@
$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}$ (13)
Inspired by this, we explore a mask structure leveraging the parity-check constraints. Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as 2nk2n-k and $d_{l}$ as nkn-k. The theoretical foundation for setting $d_{l}\approx n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k. We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$ (14)
@@
$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}$ (15)
@@
$\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$ (16)
@@ 5 Output: $(-\infty)\cdot(\neg\mathrm{~\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2. It should be noted that, in Application-Specific Integrated Circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency.
@@ 3.3 Architecture and Training Methodology
The architecture of the proposed unified error correction code Transformer is illustrated in Figure 9. The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head shared unified-attention module. This modification not only decreases computational complexity but also empowers the model to discern distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture. We employ fine-tuning techniques to enhance the model’s generalization to unseen codewords. Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
@@
$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y))$ (17)
@@
$\widetilde{\textbf{{y}}}=\left[~\left[\lvert\textbf{{y}}\rvert,~\textbf{0}_{c}\right],~\left[s(\textbf{{y}}),~\textbf{0}_{s}\right]~\right]$ (18)
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for mixed training in a unified manner. In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention.
In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each. Samples for each minibatch were randomly selected from a pool of systematic Polar [33], LDPC, and BCH codewords. The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup  [34], gradually reducing it to $10^{-6}$ by the end of the training process. We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures  [35]. The AWGN noise introduced to the links had its Signal-to-Noise Ratio (SNR) randomly chosen between 3 and 7 for each batch. All experiments were conducted on NVIDIA GeForce RTX 4090 24GB GPUs.
@@ 4 Experiments
We conducted experiments on linear block codes, specifically Polar, LDPC, and BCH, to assess the effectiveness of the proposed architecture. The parity-check matrices utilized in the paper were sourced from  [36]. We benchmark our approach against the most recent state-of-the-art performance from  [24], which trains each codeword parameter independently, and the hybrid-trained FECCT  [25]. Additionally, we compare it to legacy BP-based decoders, including unlearned BP  [37], fully supervised Hyper BP  [38], and Autoregressive BP (AR BP)  [39]. The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e. $E_{b}/N_{0}$) values, with at least $10^{5}$ random codewords tested per code type. We define aggregate BER (ABER) and aggregate BLER (ABLER) as the overall error rates spanning all evaluated codewords.
In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were directly obtained from their respective papers  [24, 25], we fine-tuned the pre-trained UECCT models on Consultative Committee for Space Data Systems (CCSDS)  [40] and MacKay  [41] codewords leveraging the hybrid-trained Polar, LDPC, and BCH models. Given GPU memory constraints and the fact that the performance of long codes has been thoroughly explored  [42], while the error correction of short codes remains far from the Shannon limit  [43], this work primarily focuses on training and validating medium-to-short codes.
The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the number of unified attention heads $H$, the dimensions of word embeddings $d_{k}$, the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$, and the dimension $d_{f}$ of the FFN layer. We present the performance of our framework with hyperparameters are set to $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=4\cdot H\cdot d_{k}$. The value of $d_{l}=64$ corresponds to the maximum syndrome length among the trained codewords.
@@
Table 1: A comparison of the negative natural logarithm of the BER
Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $E_{b}/N_{0}$ values of 4, 5, and 6 dB. It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold. Additional BER plots for Polar, LDPC and BCH codes are provided in Figure 10, 11 and Figure 12, respectively. Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT  [24] and the hybrid-trained FECCT  [25] across the majority of codewords, validating the effectiveness of the proposed architecture. Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the efficiency and power of our design. The superior performance stems from its ability to leverage a shared attention mechanism across heads, as detailed in Section 3, combined with sparsity-aware masking that reduces computational complexity while preserving decoding accuracy.
@@ 5 Analysis
We investigate the impact of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and the effectiveness of sparse masked attention. Additionally, we compare the computational complexity of our approach with existing methods.
@@ 5.1 Impact of Encoder Layers and Embedding Dimensions
Deeper architectures better capture intricate data patterns and dependencies, crucial for error correction. Additional layers enable multi-level abstraction and refinement, improving error interpretation and correction, especially in complex wireless communication systems where simpler models falter.

### v4

@@ 3.1 Shared Unified Attention
The vanilla self-attention mechanism in Transformer architectures establishes dense correlations among all input features within a sequence, resulting in a computational complexity of $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$, where $N$ is the sequence length and $d_{k}$ is the embedding dimension. This approach, while effective for capturing long-range dependencies, is computationally intensive and overlooks the intrinsic relationships across different samples, particularly in channel decoding tasks involving diverse linear block codes. Moreover, prior studies suggest that the input to the softmax function in self-attention is probabilistically sparse, with only a few weights dominating the attention distribution [28]. This sparsity implies that a low-rank approximation could effectively capture the dominant correlations, reducing redundancy without compromising performance.
@@
To investigate this hypothesis, we trained an ECCT model [22] with parameters $L=6$ (number of Transformer encoder layers), $H=8$, $d_{k}=\text{max}(2n-k)=182$, and $d_{f}=2912$ on a diverse set of linear block codes, including LDPC(49,24), LDPC(121,60), LDPC(121,80), Polar(32,16), Polar(64,32), Polar(128,86), BCH(31,16), BCH(63,36), and BCH(127,120). We analyzed the rank distribution of the $\textbf{{Q}}\cdot\textbf{{K}}^{T}$ matrix across 1000 batches for the six-layer Transformer encoder. As illustrated in Figure 5, the rank ranges from 12 to 181, with over 90% of values below 69, consistently indicating a low-rank structure across all layers and code types. This observation suggests that the high-dimensional self-attention matrix can be approximated with a lower-rank representation, offering a pathway to reduce computational overhead.
@@
Furthermore, the conventional multi-head self-attention mechanism computes independent attention scores for each head using distinct weight matrices $\textbf{{W}}_{i}^{Q}$, $\textbf{{W}}_{i}^{K}$, and $\textbf{{W}}_{i}^{V}$. However, we hypothesize that the attention distributions across heads may exhibit high similarity, particularly in channel decoding, where the sparse structure of the parity-check matrix $\mathbf{H}$ constrains bit interactions. To measure attention head similarity, we employed the Jensen-Shannon Divergence (JSD) [29], a symmetric measure of similarity between probability distributions, defined as:
$\text{JSD}(\mathbf{A}_{i}[k]\|\mathbf{A}_{j}[k])=\frac{1}{2}\left[D_{KL}(\mathbf{A}_{i}[k]\|\mathbf{M}[k])+D_{KL}(\mathbf{A}_{j}[k]\|\mathbf{M}[k])\right],$ (9)
@@
$\text{UniAttn}\left({{\textbf{{A}}_{l}},{\textbf{{V}}_{l}}}\right)=\text{Softmax}\left({{\textbf{{A}}_{l}}}\right)\cdot\left({{\textbf{{V}}_{l}^{T}}\textbf{{X}}}\right),$ (10)
@@
$\text{Attn}\left(\textbf{{W}}^{\textit{Q}},\textbf{{W}}^{\textit{K}},\textbf{{W}}^{\textit{V}}\right)=\text{Softmax}\left(\frac{\left(\textbf{{XW}}^{\textit{Q}}\right)\cdot{\left(\textbf{{XW}}^{\textit{K}}\right)}^{T}}{\sqrt{d_{k}}}\right)\cdot\left(\textbf{{XW}}^{\textit{V}}\right).$ (11)
By eliminating the projections $\textbf{{W}}^{\textit{Q}}$ and $\textbf{{W}}^{\textit{K}}$, the UAM reduces the computational complexity from $\mathcal{O}(N^{2}\cdot d_{k}+N\cdot d_{k}^{2})$ to $\mathcal{O}(N\cdot d_{l}\cdot d_{k})$, where $d_{l}\ll N$. The shared attention matrix $\mathbf{A}_{l}$ is applied across all heads, leveraging the observed head similarity to promote a unified decoding architecture. This design not only reduces the parameter count but also enables the model to learn shared features across diverse code types, facilitating a code-agnostic framework.
@@ 3.2 Sparse Masked Unified Attention
$\text{UniAttn}\left(\boldsymbol{A}_{l},\boldsymbol{V}_{l}\right)=\text{Softmax}\left(\boldsymbol{A}_{l}+\text{M}(\textbf{H})\right)\cdot\left(\boldsymbol{V}_{l}^{T}\boldsymbol{X}\right).$ (12)
@@
$\left[\overline{\textbf{{y}}},\text{s}\left(\textbf{{y}}\right)\right]\cdot\overline{\textbf{H}}=\textbf{0}.$ (13)
Inspired by this, we explore a mask structure leveraging the parity-check constraints. Let $\textbf{{A}}_{l}$ be a matrix in $\mathbb{R}^{N\times d_{l}}$, where $N$ is defined as 2nk2n-k and $d_{l}$ as nkn-k. The theoretical foundation for setting $d_{l}=n-k$ stems from the maximum rank of the parity-check matrix $\textbf{H}\in\mathbb{F}_{2}^{(n-k)\times n}$, which is nkn-k. We then construct the extended parity-check matrix $\overline{\textbf{H}}$ as follows:
$\overline{\textbf{H}}=\begin{bmatrix}{\textbf{H}}^{T}\\ {\textbf{{I}}}_{n-k}\end{bmatrix}$, (14)
@@
$H_{d}=\frac{\text{sum}(\overline{\textbf{H}})}{(2n-k)\times(n-k)}.$ (15)
@@
$\textbf{H}=\begin{bmatrix}1&1&1&0&1&0&0\\ 1&0&1&1&0&1&0\\ 0&1&1&1&0&0&1\end{bmatrix}$. (16)
@@ 5 Output: $(-\infty)\cdot(\neg\mathrm{~\overline{\textbf{H}}})$
This illustration demonstrates that the application of the mask results in sparser attention, a phenomenon particularly pronounced in low-density parity-check codes. Most importantly, the computational complexity of the UAM is now reduced to the density of the code $\mathcal{O}(N\times d_{l}\times d_{k}\times H_{d})$. Our experimental results indicate that the use of this mask has successfully reduced the computational complexity of the attention mechanism by an average of 86%. Furthermore, we provide a summary of the Python-style sparse mask construction within Algorithm 2. It should be noted that, in application-specific integrated circuit (ASIC) implementations, computations involving the masked parts can be directly skipped to enhance efficiency [31].
@@ 3.3 Architecture and Training Methodology
The architecture of the proposed unified error correction code Transformer is illustrated in Figure 9. The primary innovation in our model involves substituting the vanilla attention mechanism with a multi-head shared unified-attention module. This modification not only decreases computational complexity but also empowers the model to identify distinctive features among diverse codewords, thereby facilitating the establishment of a unified decoding architecture. We employ fine-tuning techniques to enhance the model’s generalization to unseen codewords. Furthermore, we have designed standardization units to standardize the lengths of codewords and syndromes across diverse datasets, ensuring the model’s compatibility with various code types, lengths, and rates.
@@
$\text{loss}=-\sum_{i=1}^{n}{\tilde{z}}_{i}\text{log}(f_{\theta}(y))+(1-{\tilde{z}}_{i})\text{log}(1-f_{\theta}(y)).$ (17)
@@
$\widetilde{\textbf{{y}}}=\left[~\left[\lvert\textbf{{y}}\rvert,~\textbf{0}_{c}\right],~\left[s(\textbf{{y}}),~\textbf{0}_{s}\right]~\right],$ (18)
where $\textbf{0}_{c}$ and $\textbf{0}_{s}$ represent zero vectors with sizes $N_{\text{max}}-\text{size}(\textbf{{y}})$ and $S_{\text{max}}-\text{size}(s(\textbf{{y}}))$, respectively. Through this unit, the Transformer neural decoder can accommodate any codeword within the maximum length, allowing for joint training in a unified manner. In ASIC implementations, computations for zero-padded parts can be skipped to improve efficiency, similar to the sparse-masked attention [31].
In the training phase, we conducted 1000 epochs, each consisting of 1000 minibatches, which in turn contained 512 samples each. Samples for each minibatch were randomly selected from a pool of LDPC, Polar, and BCH codewords. The learning rate was initialized at $10^{-3}$, and a cosine decay scheduler was applied without warmup [32], gradually reducing it to $10^{-6}$ by the end of the training process. We employ the Adam optimizer to dynamically adjust the learning rate, leveraging its efficacy in managing diverse data types and model architectures [33]. The AWGN noise introduced to the links had its signal-to-noise ratio (SNR) randomly chosen between 3 and 7 for each batch. All experiments were conducted on NVIDIA GeForce RTX 4090 24GB GPUs.
@@ 4 Experiments
We conducted experiments on linear block codes, specifically LDPC, Polar, and BCH, to assess the effectiveness of the proposed architecture. The parity-check matrices utilized in the paper were sourced from [34]. We benchmark our approach against the most recent state-of-the-art methods from [22], which trains each codeword parameter independently, and the jointly trained FECCT [23]. Additionally, we compare it to legacy belief propagation (BP)-based decoders, including unlearned BP [35], fully supervised Hyper BP [36], and Autoregressive BP (AR BP) [37]. The results are presented in the form of bit error rate (BER) and block error rate (BLER) at various normalized SNR (i.e. $\mathrm{E_{b}/N_{0}}$) values, with at least $10^{5}$ random codewords tested per SNR. We define aggregate BER (ABER) and aggregate BLER (ABLER) as the overall error rates spanning all evaluated codewords.
In addition, to enable direct comparison with ECCT and FECCT, whose simulation results were sourced from their respective papers [22, 23], we fine-tuned the pre-trained UECCT models on consultative committee for space data systems (CCSDS) [38] and MacKay [39] codewords using jointly trained LDPC, Polar, and BCH models. Given GPU memory constraints and the fact that the performance of long codes has been thoroughly explored [40], while the error correction of short codes remains far from the Shannon limit [41], this work primarily focuses on training and validating medium-to-short codes.
The proposed unified error correction code Transformer is defined by the following hyperparameters: the number of Transformer encoder layers $L$, the number of unified attention heads $H$, the dimensions of word embeddings $d_{k}$, the dimensions $d_{l}$ of the trainable memory $\textbf{{A}}_{l}$ and $\textbf{{V}}_{l}$, and the dimension $d_{f}$ of the FFN layer. We present the performance of our framework with hyperparameters are set to $L=6$, $H=8$, $d_{k}=64$, $d_{l}=64$, and $d_{f}=4\cdot H\cdot d_{k}$. The value of $d_{l}=64$ corresponds to the maximum syndrome length among the jointly trained codewords.
Table 1: Negative natural logarithm of BER for ECCT, FECCT, and UECCT across Polar, LDPC, BCH, CCSDS, and MacKay codes at $\mathrm{E_{b}/N_{0}}$ = 4, 5, 6 dB, with best results in bold (higher is better).
@@
Table 1 presents the simulation results, specifically showing the negative natural logarithm of the BER at $\mathrm{E_{b}/N_{0}}$ values of 4, 5, and 6 dB. It is noted that higher values in this context indicate better performance, with the best results for each SNR highlighted in bold. Additional BER plots for Polar, LDPC and BCH codes are provided in Figure 10, 11 and Figure 12, respectively. Simulation results demonstrate that our proposed UECCT outperforms both the individually trained ECCT [22] and the jointly trained FECCT [23] across the majority of codewords, validating the effectiveness of the proposed architecture. Notably, UECCT even surpasses FECCT’s performance with an embedding length of 128 when operating at a more compact embedding length of 64, further highlighting the computational efficiency of our design. The superior performance arises from its ability to utilize a shared attention mechanism across heads, combined with sparsity-aware masking that reduces computational complexity while maintaining decoding accuracy.
Figure 13: BER comparison between the NMS algorithm and the proposed UECCT for medium-to-long LDPC codes.
To further validate the performance and scalability of the proposed method on medium-to-long codes, we conducted joint training using LDPC(576, 288), LDPC(768, 512), and LDPC(960, 800) codewords with hyperparameters set to $L=12$, $H=8$, $d_{k}=16$, $d_{l}=288$, and $d_{f}=512$. The neural network training and inference were performed under an AWGN channel with BPSK modulation. During training, the SNR was randomly selected between 2 dB and 3 dB, with a batch size of 192, while other training methods were consistent with those described in Section 3.3. During inference, $10^{5}$ code blocks were tested per SNR. For comparison with traditional decoding algorithms, we utilized the normalized min-sum (NMS) algorithm for LDPC codes [42]. The NMS algorithm was configured with a normalization factor of 0.875 and a maximum of 15 iterations. Additionally, an early stopping mechanism was implemented to halt decoding once the parity-check matrix $\mathbf{H}$ is satisfied. Figure 13 presents the BER performance comparison between the NMS algorithm and the proposed UECCT. The simulation results demonstrate that the proposed method achieves performance comparable to the traditional NMS decoding algorithm. This indicates that our decoding architecture effectively scales to medium-to-long codes.
@@ 5 Analysis
Using the parameters in Table 1, we examine the effect of Transformer encoder layers on performance, focusing on the embedding dimension $d_{k}$ and sparse masked attention efficiency. We also compare our approach’s computational complexity with that of existing methods.
@@ 5.1 Impact of Encoder Layers and Embedding Dimensions
Deeper architectures better capture intricate data patterns and dependencies, crucial for error correction. Additional layers enable multi-level abstraction and refinement, improving error estimation and correction, especially in complex wireless communication systems where simpler models falter.